OpenStudy (anonymous):

What are the vertex and the axis of symmetry of the equation? y = –2x^2 + 8x – 18 (1 point) vertex: ( 2, –10) axis of symmetry: x = 2 vertex: (2, –10) axis of symmetry: x = –10 vertex: (–2, –10) axis of symmetry: x = –2 vertex: (–2, 10) axis of symmetry: y = –2

OpenStudy (anonymous):

The graph below models the acceleration of a car. Interpret the meaning of the vertex. The car traveled 10 miles in 60 seconds. It took the car 20 seconds to accelerate. It took the car 10 seconds to reach a speed of 60 miles per hour. It took the car 20 seconds to reach a speed of 60 miles per hour.

OpenStudy (anonymous):

y= -2(x^2-4x+9)=-2( x-2)^2 -18+8 (y+10) = -2( x-2)^2

OpenStudy (anonymous):

does it help u for 1st Q?

OpenStudy (anonymous):

vertex= (2,-10) axis of sym x=2

OpenStudy (dayianaforever):

what was the answer